### v1

# Central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discretely valued fields of dimension 2

Abstract
Let $K$ be a complete discretely valued field with residue field $\bar{K}$ of dimension $1$ (not necessarily perfect). This occurs if and only if $K$ has dimension $2$. We prove the following statements on the arithmetic of such fields:
  • The “period equals index” property holds for central simple $K$-algebras.
  • For every prime $p$, every class in the Milnor $\mathrm{K}$-theory modulo $p$ is represented by a symbol.
  • Serre’s Conjecture II holds for the field $K$. That is, for every semisimple and simply connected $K$-group $G$, the set $H^{1}(K,G)$ is trivial.
MSC Classes: 16K50, 19F15, 11E72, 12G05, 12G10, 20G10.
Keywords: Period-index problem, Milnor $\mathrm{K}$-theory, Galois cohomology, principal homogeneous spaces, Serre’s conjecture II.
1.   Introduction
In this article, we investigate various classical algebraic and arithmetic questions related to central simple algebras, Milnor $K$-theory and homogeneous spaces over complete discrete valuation fields with dimension $2$. Most of the results we provide are well-known when the residue field is assumed to be perfect, and our goal is remove that assumption.
Central simple algebras and period-index problem.
The period-index problem is a classic in the theory of Brauer groups of fields. The period $\mathrm{per}(\alpha)$ of a Brauer class $\alpha$ is its order in the Brauer group and its index $\mathrm{ind}(\alpha)$ is the degree of the division algebra underlying a representative of the class. It is well-known that $\mathrm{per}(\alpha)|\mathrm{ind}(\alpha)$ and that both have the same prime factors (cf. for instance [GS17, Thm. 2.8.7]). Relations of the type $\mathrm{ind}(\alpha)|\mathrm{per}(\alpha)^{n}$ are known for many fields having a nice algebraic, geometric or arithmetic behaviour (e.g. [Art87, AAIKL, Ben24, BHN32, CTOP02, dJ04, FS89, Lie15, Sal97, PS14]). It is often reasonable to expect the equality $\mathrm{ind}(\alpha)=\mathrm{per}(\alpha)$ for Brauer classes $\alpha$ over a field with dimension $2$. In §3, we use an argument due to Kato to prove that such a result holds over any complete discrete valuation field $K$ with dimension $2$ (cf. Propositions 3.1 and 3.2). In the case where $K$ has characteristic 00 and contains all $p$-th roots of unity, where $p$ is the characteristic exponent of its residue field, arguments of Saltman allow us to check a stronger result, stating that all central simple algebras over $K$ are cyclic.
Milnor $\mathrm{K}$-theory.
If $K$ is a field containing the $p$-th roots of unity for a given prime $p$, the Bloch-Kato conjecture (a result now proved by Rost and Voevodsky) provides an isomorphism
$\mathrm{K}^{\mathrm{M}}_{2}(K)/p\xrightarrow{\sim}{\rm{Br}}(K)[p],$
sending the class of the symbol $\{x,y\}$ to the class of the cyclic algebra $(x,y)_{\zeta_{p}}$. This tells us that the cyclicity of central simple algebras of period $p$ can be restated in terms of Milnor $\mathrm{K}$-theory as every class in $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ being represented by a symbol. One could expect then such a result for complete discretely valued fields of dimension 2 without the assumption on roots of unity. One of the main results of this article, proved in §4, corresponds to this statement for fields of characteristic 00:
Theorem 1.1 (Theorem 4.1).
Let $K$ be a complete discretely valued field of characteristic 00 with residue field $\bar{K}$ of dimension $1$. Let $p$ be a prime number. Then every class in $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ is represented by a symbol.
In order to tackle this statement, the main tool is a description of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic given by Kato. However, this description works under the hypothesis that $K$ contains the $p$-th roots of unity. We remove this assumption in Proposition 4.3. Theorem 1.1 readily follows by analogy with Saltman’s arguments, which we present in this setting in the appendix.
Serre’s Conjecture II
Period-index questions have important applications to the arithmetic of homogeneous spaces over low-dimensional fields, and more precisely to the so-called Serre’s Conjecture II:
Conjecture 1.2 (Serre’s conjecture II).
Let $K$ be a field of dimension $\leq 2$. Then every principal homogeneous space under a semisimple simply connected $K$-group has a rational point.
This conjecture remains open, although a lot has been done in particular cases, whether specifying the type of the group (cf. for instance [BP95, BFT07, Gar01, KMRT98, Gar10, Che03, Gil01]) or the field of definition (cf. [Kne65a, Kne65b, BT87, Har65, Har66, Che89, CTGP04, dJHS11, Ben24]. For a survey on the topic, see the lecture notes [Gil19]. More recently, in [ILA24], it is proved that Serre’s Conjecture II can be reduced to the case of countable characteristic 00 fields of dimension $2$.
In this article, we settle the following case of Serre’s Conjecture II:
Theorem 1.3 (Theorem 6.1).
Let $K$ be a complete discretely valued field of dimension $2$. Let $G$ be a semisimple simply connected $K$-group. Then $H^{1}(K,G)=1$.
When the residue field $\bar{K}$ is perfect, this result was already known by work of Kneser in the p-adic case [Kne65a, Kne65b] and Bruhat and Tits [BT87, Cor. 3.15]. The cases of groups of classical type and of types $F_{4}$ and $G_{2}$ were also already known by the aforementioned result of Bayer and Parimala [BP95]. In §6, we deduce the cases of trialitarian $D_{4}$, $E_{6}$ and $E_{7}$ from Propositions 3.1 and 3.2, while we obtain the case of $E_{8}$ by two methods: when $K$ has characteristic 00, we study the maximal solvable extension of $K$ of degree $2^{a}3^{b}5^{c}$ (cf. Proposition 6.2) and use an argument due to the first author; when $K$ has positive characteristic, we follow the original arguments by Bruhat and Tits in [BT87] and adapt them to the case of imperfect residue fields.
Outline of the article
In §2, we give some preliminaries on the various existing notions of dimension for fields, on central simple algebras and on Milnor $\mathrm{K}$-theory, recalling in particular an explicit description given by Kato of the $\mathrm{K}$-theory modulo $p$ for complete discretely valued fields of mixed characteristic with imperfect residue field. In §3, we settle the “period equals index” property for complete discretely valeud fields of dimension 2 in arbitrary characteristic. In §4, we work in mixed characteristic, we remove the hypothesis on the roots of unity in Kato’s description mentioned before and we apply it in order to prove Theorem 1.1. In order to do so, we state three “absorption lemmas” that help us to reduce sums of symbols to a single symbol (Lemmas 4.4, 4.5 and 4.6). The proof of these lemmas is completely analogous to computations previously carried by Saltman and thus are given in an appendix to this article. In §5, we consider the case where $K$ has positive characteristic and give some partial results in this direction. Finally, in §6, we prove Theorem 1.3 separating the cases of characteristic 00 and positive characteristic.
Acknowledgements
The authors would like to thank Skip Garibaldi for his suggestions that allowed to improve the exposition, as well as Raman Parimala and Suresh Venapally for their interest. The first author was supported by the project “Group schemes, root systems, and related representations” founded by the European Union - NextGenerationEU through Romania’s National Recovery and Resilience Plan (PNRR) call no. PNRR-III-C9-2023-I8, Project CF159/31.07.2023, and coordinated by the Ministry of Research, Innovation and Digitalization (MCID) of Romania. The third author’s research was partially supported by ANID via FONDECYT Grant 1240001.
2.   Notations and Preliminaries
The following notations concerning complete discretely valued fields will be used throughout the whole article, except in sections §§2.12.3, where we recall some results over arbitrary fields:
  • $K$ stands for a field that is complete with respect to a discrete valuation $v_{K}$;
  • $\mathcal{O}_{K}$ stands for its integer ring and $\bar{K}$ for its residue field, which can be arbitrary;
  • for an element $x\in\mathcal{O}_{K}$, its image in $\bar{K}$ is denoted $\bar{x}$.
We start by recalling some basic facts about the dimension of fields, central simple algebras and Milnor $\mathrm{K}$-theory.
2.1 The dimension of a field
The cohomological dimension $\mathrm{cd}(K)$ of a field $K$ is the cohomological dimension of its absolute Galois group. In other words, it is the smallest integer $\delta$ (or $\infty$ if such an integer does not exist) such that $H^{n}(K,M)=0$ for every $n>\delta$ and each finite Galois module $M$. This notion turns out not to be completely satisfactory for imperfect fields, in particular because a field with characteristic $p>0$ and whose Galois group is a pro-$p$-group has cohomological dimension at most $1$. Better definitions in that context were provided by Kato in [Kat82] and by Gille in [Gil00].
Definition 2.1.
Let $K$ be any field and $p$ a prime number.
  • (i)
    Assume that the characteristic of $K$ is different from $p$. The $p$-dimension $\mathrm{dim}_{p}(K)$ and the separable $p$-dimension $\mathrm{scd}_{p}(K)$ of $K$ are both equal to the $p$-cohomological dimension of the absolute Galois group of $K$.
  • (ii)
    Assume now that $K$ has characteristic $p$. Let $\Omega^{i}_{K}$ be the $i$-th exterior product over $K$ of the absolute differential module $\Omega^{1}_{K/\mathbb{Z}}$ and let $H^{i+1}_{p}(K)$ be the cokernel of the morphism $\mathfrak{p}^{i}_{K}\,:\,\Omega^{i}_{K}\to\Omega^{i}_{K}/d(\Omega^{i-1}_{K})$ defined by
    $x\frac{dy_{1}}{y_{1}}\wedge...\wedge\frac{dy_{i}}{y_{i}}\mapsto(x^{p}-x)\frac{dy_{1}}{y_{1}}\wedge...\wedge\frac{dy_{i}}{y_{i}}\mod d(\Omega^{i-1}_{K}),$
    for $x\in K$ and $y_{1},...,y_{i}\in K^{\times}$. The $p$-dimension $\mathrm{dim}_{p}(K)$ of $K$ is the smallest integer $\delta$ (or $\infty$ if such an integer does not exist) such that $[K:K^{p}]\leq p^{\delta}$ and $H^{\delta+1}_{p}(L)=0$ for all finite extensions $L$ of $K$. The separable $p$-dimension $\mathrm{scd}_{p}(K)$ of $K$ is the smallest integer $i$ (or $\infty$ if such an integer does not exist) such that $H^{i+1}_{p}(L)=0$ for all finite separable extensions $L$ of $K$.
  • (iii)
    The dimension $\mathrm{dim}(K)$ of $K$ is the supremum of all the $\mathrm{dim}_{\ell}(K)$’s when $\ell$ runs through all prime numbers. The separable dimension $\mathrm{scd}(K)$ of $K$ is the supremum of all the $\mathrm{scd}_{\ell}(K)$’s when $\ell$ runs through all prime numbers.
Note that, for any prime number $\ell$, we have $\mathrm{scd}_{\ell}(K)\leq\dim_{\ell}(K)$, and hence $\mathrm{scd}(K)\leq\dim(K)$.
For the reader’s convenience, we briefly recall the behaviour of these different notions of dimension.
Lemma 2.2.
Let $K$ be a field.
  1. (i)
    A field $K$ has dimension 00 if, and only if, $K$ is algebraically closed.
  2. (ii)
    Let $K$ be a field with $[K:K^{p}]\leq p$ where $p$ is the characteristic exponent. Then $K$ has dimension at most $1$ if, and only if, ${\rm{Br}}(L)=0$ for each finite extension $L/K$.
Proof.
We denote by $p$ the characteristic exponent of $K$.
  1. (i)
    If $K$ is algebraically closed, then it is straightforward that $\mathrm{dim}(K)=0$. Conversely assume that $\mathrm{dim}(K)=0$. Then $\mathrm{dim}_{\ell}(K)=0$ for each $\ell\neq p$, and so the $-\ell$-Sylow subgroups of $\mathop{\rm Gal}\nolimits(K^{\mathrm{sep}}/K)$ are trivial. It follows that $\mathop{\rm Gal}\nolimits(K^{\mathrm{sep}}/K)$ is a pro-$p$-group. The case $p=1$ is thus finished. Assuming $p>1$, we have $[K:K^{p}]=1$ so that $K$ is perfect. Furthermore we have
    $0=H^{1}_{p}(K)=K/\{x^{p}-x\mid x\in K\}=H^{1}(\mathop{\rm Gal}\nolimits(K^{\mathrm{sep}}/K),{\mathbb{Z}}/p{\mathbb{Z}}),$
    so that $\mathop{\rm Gal}\nolimits(K^{\mathrm{sep}}/K)=1$ (see [Ser02, I, §.4.1, Prop. 21]). The field $K$ is perfect and is separably closed hence is algebraically closed.
  2. (ii)
    According to [GS17, Thm. 6.1.8], for $\ell\neq p$, the field $K$ has $\ell$-dimension at most $1$ if, and only if, ${\rm{Br}}(L)[\ell]=0$ for each finite extension $L/K$. It remains to check the same statement when $\ell=p>1$. This follows from the identification ${\rm{Br}}(L)[p]=H^{2}_{p}(L)$, cf [GS17, Thm. 9.2.4].
Example 2.3.
a
  1. (a)
    Assume that $K$ is a $C_{1}$ field, that is, a field over which any projective hypersurface in $\mathbb{P}^{n}_{K}$ of degree $d\leq n$ has a rational point. If $K$ is not algebraically closed, then $\mathrm{dim}(K)=1$. Indeed, according to [GS17, Prop. 6.2.3], we have $\dim_{\ell}(K)\leq 1$ for each $\ell\not=p$ and also ${\rm{Br}}(L)=0$ for each finite field extension $L$ of $K$. This shows that $\mathrm{dim}(K)\leq 1$ when $p=1$. Now, if $p>1$, we have $H^{2}_{p}(L)={\rm{Br}}(L)[p]=0$ for each finite field extension $L$ of $K$ by [GS17, Thm. 9.2.4] and $[K:K^{p}]\leq p$ by [Ser02, II, §3.2, Cor. to Prop. 8], so that $\mathrm{dim}_{p}(K)\leq 1$. Since $K$ is not algebraically closed, we conclude that $\mathrm{dim}(K)=1$.
  2. (b)
    By taking into account respectively theorems by Chevalley, Tsen and Lang (see [Ser02, II, §3.3]), we have $\mathrm{dim}(K)=1$ in the following cases:
    • $K$ is a finite field;
    • $K=K_{0}(C)$ is the function field of an algebraic curve $C$ over an algebraically closed field $K_{0}$;
    • $K$ is a henselian discrete valued field such that its completion $\widehat{K}$ is separable over $K$ and whose residue field $\bar{K}$ is algebraically closed.
  3. (c)
    There are more examples of fields of dimension $\leq 1$, see for example the field constructed by Colliot-Thélène and Madore [CTM04, Thm. 1.1].
  4. (d)
    Assume that $K$ is separably closed of characteristic $p\geq 2$ and that $[K:K^{p}]=p$. Then $\mathrm{dim}(K)=1$. Indeed each finite field extension $L$ of $K$ is separably closed so that ${\rm{Br}}(L)=0$.
  5. (e)
    A very important feature of Definition 2.1 is that, for any complete discretely valued field $K$ with residue field $\bar{K}$, we have $\dim(K)=\dim(\bar{K})+1$, cf. [Kat82, Cor. to Thm. 3].
2.2 Central simple algebras
Let $K$ be a field. According to Wedderburn’s Theorem, a central simple algebra over $K$ is always of the form $\mathcal{M}_{n}(D)$ for some $n\geq 1$ and some central division algebra $D$. Two central simple algebras are said to be Brauer equivalent if the underlying central division algebras are isomorphic. Central simple algebras up to Brauer equivalence form a group with respect to the tensor product, called the Brauer group of $K$ and denoted ${\rm{Br}}(K)$. The Brauer group ${\rm{Br}}(K)$ coincides with the Galois cohomology group $H^{2}(K,\mathbb{G}_{\mathrm{m}})$ provided that $K$ is perfect. All these considerations allow to define two invariants associated to a given central simple algebra $A$ over $K$:
  • the period of $A$ is the order of its class in ${\rm{Br}}(K)$;
  • if $D$ is the underlying central division algebra of $A$, the dimension of $D$ is a perfect square and its square root is called the index of $A$.
As one can observe, both notions are actually well-defined for the class of $A$ in the Brauer group, so that one can talk about period and index of a Brauer class instead of a central simple algebra.
[…]
$\langle x,y:x^{m}=a,y^{m}=b,xy=\zeta yx\rangle.$
One often writes $(a,b)_{\zeta}$ instead of $(\chi,b)$.
Central simple algebras of the form $(\chi,b)$ are usually called cyclic algebras, and period is always equal to index for such algebras.
2.3 Milnor $\mathrm{K}$-theory
Let $K$ be any field. The Milnor $\mathrm{K}$-theory groups of $K$ are defined and studied in full generality in [GS17, Ch. 7] (and unless otherwise stated, every statement presented in this subsection comes from there). In this article, we will only use the first three Milnor $\mathrm{K}$-theory groups, which can be explicitly described as:
$\mathrm{K}^{\mathrm{M}}_{0}(K):=\mathbb{Z},\quad\mathrm{K}^{\mathrm{M}}_{1}(K):=K^{\times}$
$\mathrm{K}^{\mathrm{M}}_{2}(K):=(K^{\times}\otimes_{\mathbb{Z}}K^{\times})/\left\langle x\otimes(1-x):x\in K\smallsetminus\{0,1\}\right\rangle.$
For $x,y\in K^{\times}$, the symbol $\{x,y\}$ denotes the class of $x\otimes y$ in $\mathrm{K}^{\mathrm{M}}_{2}(K)$.
When $L$ is a finite extension of $K$ and $q\in\{0,1,2\}$, we have a norm homomorphism
$N_{L/K}:\mathrm{K}^{\mathrm{M}}_{q}(L)\rightarrow\mathrm{K}^{\mathrm{M}}_{q}(K),$
satisfying the following properties :
  • $\bullet$
    For $q=0$, the map $N_{L/K}:\mathrm{K}^{\mathrm{M}}_{0}(L)\rightarrow\mathrm{K}^{\mathrm{M}}_{0}(K)$ is given by multiplication by $[L:K]$.
  • $\bullet$
    For $q=1$, the map $N_{L/K}:\mathrm{K}^{\mathrm{M}}_{1}(L)\rightarrow\mathrm{K}^{\mathrm{M}}_{1}(K)$ coincides with the usual norm $L^{\times}\rightarrow K^{\times}$.
  • $\bullet$
    The equality $N_{L/K}(\{x,y\})=\{x,N_{L/K}(y)\}$ holds for $x\in\mathrm{K}^{\mathrm{M}}_{1}(K)$ and $y\in\mathrm{K}^{\mathrm{M}}_{1}(L)$.
Milnor $\mathrm{K}$-theory is also endowed with residue maps. Indeed, when $K$ is a complete discretely valued field with ring of integers $\mathcal{O}_{K}$ and residue field $\bar{K}$, there exists a unique residue morphism:
$\partial:\mathrm{K}^{\mathrm{M}}_{2}(K)\rightarrow\mathrm{K}^{\mathrm{M}}_{1}(\bar{K}),$
such that, for any uniformizer $\pi$ and for each unit $u\in\mathcal{O}_{K}^{\times}$, one has:
($\partial(\{\pi,u\})=\bar{u}.$
The kernel of $\partial$ is the subgroup $U_{2}(K)$ of $\mathrm{K}^{\mathrm{M}}_{2}(K)$ generated by symbols of the form $\{x,y\}$ with $x,y\in\mathcal{O}_{K}^{\times}$. We denote by $U_{2}^{i}(K)$ the subgroup of $\mathrm{K}^{\mathrm{M}}_{2}(K)$ generated by those symbols that lie in $U_{2}(K)$ and that are of the form $\{x,y\}$ with $x\in 1+\pi^{i}\mathcal{O}_{K}$ and $y\in K^{\times}$.
The residue map turns out to be very useful to understand the group $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ when $p$ is a prime number different from the characteristic of $\bar{K}$. When $p$ is equal to $\mathrm{char}(\bar{K})$ and $K$ contains a primitive $p$-th root of unity $\zeta_{p}$, one can use instead a precise description of $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ due to Kato. For that purpose, we set $\mathrm{k}^{\mathrm{M}}_{2}(K):=\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ and $u_{2}^{i}(K):=U_{2}^{i}(K)/p$, and we define $\epsilon_{K}:=\frac{v_{K}(p)p}{p-1}$. Then, given a uniformizer $\pi\in K$, we have the following isomorphisms (cf. [Kat82, Thm. 2]):
$\rho^{2}_{0}:\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})\oplus\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})$ $\to\mathrm{k}^{\mathrm{M}}_{2}(K)/u^{1}_{2}(K),$ (1)
($\left(\{\bar{x},\bar{y}\},0\right)$$\mapsto\{x,y\},$
$\left(0,\{\bar{y}\}\right)$$\mapsto\{\pi,y\};$
for every $0<i<\epsilon_{K}$ such that $(i,p)=1$ and any $\omega$ with $v_{K}(\omega)=i$,
$\rho^{2}_{i}:\Omega^{1}_{\bar{K}}$ $\to u^{i}_{2}(K)/u^{i+1}_{2}(K),$ (2)
$\bar{x}\,\frac{d\bar{y}}{\bar{y}}$$\mapsto\{1+\omega x,y\};$
for every $0<i<\epsilon_{K}$ such that $p|i$ and any $\varpi$ with $v_{K}(\varpi)=i/p$;
$\rho^{2}_{i}:\Omega^{1}_{\bar{K}}/\mathrm{Ker}(d)\oplus\Omega^{0}_{\bar{K}}/\mathrm{Ker}(d)$ $\to u^{i}_{2}(K)/u^{i+1}_{2}(K),$ (3)
$\left(\bar{x}\,\frac{d\bar{y}}{\bar{y}},0\right)$$\mapsto\{1+\varpi^{p}x,y\},$
$\left(0,\bar{x}\right)$$\mapsto\{\pi,1+\varpi^{p}x\};$
for $i=\epsilon_{K}$,
$\rho^{2}_{i}:H_{p}^{2}(\bar{K})\oplus H_{p}^{1}(\bar{K})$ $\to u^{i}_{2}(K),$ (4)
$\left(\bar{x}\,\frac{d\bar{y}}{\bar{y}},0\right)$$\mapsto\{1+(\zeta_{p}-1)^{p}x,y\},$
$\left(0,\bar{x}\right)$$\mapsto\{\pi,1+(\zeta_{p}-1)^{p}x\}.$
3.   Period and index of central simple algebras
In this section, we study central simple algebras over complete discretely valued fields of dimension $2$. The following proposition can be easily derived from results due to Saltman (private notes exposed partially in [LPS]):
Proposition 3.1.
Let $K$ be a complete discretely valued field of characteristic 00 with residue field $\bar{K}$ of characteristic exponent $p$ and of dimension $1$. If $\zeta_{p}\in K$, then every central simple algebra over $K$ with prime period is cyclic. In particular, period equals index for every central simple $K$-algebra.
Proof.
We denote by $K^{\mathrm{unr}}$ the maximal unramified extension of $K$. Let $\ell$ be a prime number and take $\alpha\in{\rm{Br}}(K)[\ell]$. We need to prove that $\alpha$ can be represented by a cyclic algebra.
By [Ser79, XII, §3, Exer. 3] we have an exact sequence:
$0\to{\rm{Br}}(\bar{K})\{\ell\}\to{\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}\to H^{1}(\bar{K},{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})\to 0$ (5)
Moreover, if $\ell\neq p$, then ${\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}={\rm{Br}}(K)\{\ell\}$. Now ${\rm{Br}}(\bar{K})=0$ because $\dim(\bar{K})=1$. We therefore obtain a residue isomorphism ${\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}\cong H^{1}(\bar{K},{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})$ that shows that any element in ${\rm{Br}}(K^{\mathrm{unr}}/K)\{\ell\}$ is of the form $(\chi,\pi)$ for an arbitrary uniformizer $\pi\in K$ and some unramified character $\chi\in H^{1}(K,{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})$. This proves the claim if $\ell\neq p$. If $\ell=p$, according to the proof of [LPS, Prop. 1.8], there is a uniformizer $\pi$ of $K$ and an element $a\in K$ such that $\alpha-(a,\pi)_{\zeta_{p}}\in{\rm{Br}}(K^{\mathrm{unr}}/K)$. In particular, we can write
$\alpha=(a,\pi)_{\zeta_{p}}+(\chi,\pi)=(a,\pi)_{\zeta_{p}}+(b,\pi)_{\zeta_{p}}=(ab,\pi)_{\zeta_{p}},$
for some $b\in K$ and hence $\alpha$ is represented by a cyclic algebra, as wished.
Now, the “period equals index” property for all central simple $K$-algebras follows because it can be checked by considering only algebras with $\ell$-primary period for every prime number $\ell$ and because the case of algebras with $\ell$-primary period can itself be reduced to that of algebras of period $\ell$ according to [CTGP04, Thm. 1.1.(iv)]. ∎
Thanks to work of Kato, the result stating that period equals index for central simple $K$-algebras can in fact be generalized to the positive characteristic case in the following way:
Proposition 3.2.
Let $K$ be a complete discretely valued field with residue field $\bar{K}$ of dimension $1$. Then period equals index for every central simple $K$-algebra.
Proof.
As in the proof of Proposition 3.1, we can use [CTGP04, Thm. 1.1.(iv)] to reduce to the case of a division $K$-algebra $A$ of prime period $\ell$ and we can get a residue isomorphism ${\rm{Br}}(K^{\mathrm{unr}}/K)\cong H^{1}(\bar{K},{\mathbb{Q}}/{\mathbb{Z}})$. If $A$ is split over the unramified closure $K^{\mathrm{unr}}$, it follows that $[A]$ is of the form $(\chi,\pi)$ for some unramified character $\chi\in H^{1}(K,{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})$. Thus $A$ is isomorphic to a cyclic $K$-algebra of degree $\ell$ and in particular has index $\ell$.
It remains to deal with the case when $A\otimes_{K}K^{\mathrm{unr}}$ is not split. According to [Ser79, XII, §1, Cor. to Thm. 1], this can happen only in the case where $\bar{K}$ is imperfect and, again by [Ser79, XII, §3, Exer. 3], only if $\ell=p$. Since $\dim_{p}(\bar{K})\leq 1$, we have $[\bar{K}:\bar{K}^{p}]=p$. In this case Kato has shown that $A$ has index $p$ (cf [Kat79, §4, Lemma 5]). ∎
4.   Symbols and Milnor $K$-theory
In this section, we are interested in the second degree Milnor K-theory group of a complete discretely valued field $K$ of dimension $2$. The main theorem we prove generalizes the cyclicity result in Proposition 3.1 to Milnor K-theory classes when $K$ is not anymore assumed to contain $p$-th roots of unity, and it can be stated as follows:
Theorem 4.1.
Let $K$ be a complete discretely valued field of characteristic 00 with residue field $\bar{K}$ of dimension $1$. Let $p$ be a prime number. Then every class in $\mathrm{K}^{\mathrm{M}}_{2}(K)/p$ is represented by a symbol.
The rest of this section is devoted to the proof of this theorem. Let us first deal with the easy case when $\bar{K}$ has characteristic $\neq p$. Under that assumption, since $p$ is invertible in $\bar{K}$, by [GS17, Cor. 7.1.10], we have the exact sequence:
$0\rightarrow\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})\rightarrow\mathrm{k}^{\mathrm{M}}_{2}(K)\xrightarrow{\partial}\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})\rightarrow 0,$
where the residue map $\partial:\mathrm{k}^{\mathrm{M}}_{2}(K)\rightarrow\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})$ sends a symbol of the form $\{\pi,u\}\in\mathrm{k}^{\mathrm{M}}_{2}(K)$, with $\pi$ a uniformizer in $K$ and $u$ a unit in $\mathcal{O}_{K}$, to $\overline{u}\in\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})$. According to the Bloch-Kato conjecture, the group $\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})$ is trivial because $\bar{K}$ has dimension $1$. Hence $\partial$ is an isomorphism, and every element in $\mathrm{k}^{\mathrm{M}}_{2}(K)$ is a symbol of the form $\{\pi,u\}$.
Remark 4.2.
This proof works just as well for $K$ of positive characteristic provided that it is different from $p$. We comment on the case of characteristic $p$ in §5.
Let us now move on to the case where $\mathrm{char}(\bar{K})=p$. If $K$ contains all $p$-th roots of unity, Theorem 4.1 immediately boils down to Proposition 3.1 since the Bloch-Kato conjecture provides an isomorphism $\mathrm{k}^{\mathrm{M}}_{2}(K)\cong{\rm{Br}}(K)[p]$ in which symbols correspond to cyclic algebras. We henceforth do not assume anymore that $K$ contains all $p$-th roots of unity. In that case, we a priori do not have Kato’s isomorphisms (1)–(4). So we first need to extend them to that case:
Proposition 4.3.
Let $K$ be a complete discretely valued field of characteristic 00 with residue field $\bar{K}$ of dimension $1$ and characteristic $p$. Define $\epsilon_{K}:=\frac{pv_{K}(p)}{p-1}$. Then $u_{2}^{\lfloor\epsilon_{K}\rfloor+1}(K)$ is trivial and we have the following isomorphisms for any uniformizer $\pi\in K$:
$\rho_{0}:\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})$ $\to\mathrm{k}^{\mathrm{M}}_{2}(K)/u^{1}_{2}(K),$ (6)
$\bar{y}$$\mapsto\{y,\pi\};$
for every $0<i<\epsilon_{K}$ with $(i,p)=1$ and any $\omega$ with $v_{K}(\omega)=i$,
$\rho_{i}:\Omega^{1}_{\bar{K}}$ $\to u^{i}_{2}(K)/u^{i+1}_{2}(K),$ (7)
$\bar{x}\,\frac{d\bar{y}}{\bar{y}}$$\mapsto\{1+\omega x,y\};$
for every $0<i<\epsilon_{K}$ with $p|i$ and any $\varpi$ with $v_{K}(\varpi)=i/p$;
$\rho_{i}:\Omega^{0}_{\bar{K}}/\mathrm{Ker}(d)$ $\to u^{i}_{2}(K)/u^{i+1}_{2}(K),$ (8)
$\bar{x}\pmod{\bar{K}^{p}}$$\mapsto\{1+\varpi^{p}x,\pi\};$
and for $i=\epsilon_{K}$ when $\epsilon_{K}\in{\mathbb{Z}}$,
$\rho_{i}:H_{p}^{1}(\bar{K})$ $\to u^{i}_{2}(K),$ (9)
$[\bar{x}]$$\mapsto\{1+\pi^{i}ux,\pi\},$
for some well-chosen unit $u\in\mathcal{O}_{K}^{\times}$.
Proof.
Let us start by noting that, since $\dim(\bar{K})=1$, for any $\bar{y}\in\bar{K}^{\times}\smallsetminus(\bar{K}^{\times})^{p}$ we have
$\Omega^{0}_{\bar{K}}=\bar{K},\quad\Omega^{1}_{\bar{K}}=\bar{K}\frac{d\bar{y}}{\bar{y}},\quad\text{and}\quad\Omega^{2}_{\bar{K}}=0.$
When $\zeta_{p}\in K$, the isomorphisms from the statement and the triviality of $u_{2}^{\lfloor\epsilon_{K}\rfloor+1}(K)$ are immediately given by isomorphisms (1)–(4) and the previous computation of $\Omega^{i}_{\bar{K}}$, except that we have reversed the order of the coordinates in some symbols, which amounts to a sign change.
We assume now that $\zeta_{p}\not\in K$ and we define $K^{\prime}:=K(\zeta_{p})$. Then we know that $K^{\prime}/K$ is a totally ramified cyclic extension of degree $d$ dividing $p-1$. More precisely, we know that there exists a uniformizer $\pi\in K$ such that $K^{\prime}=K(\sqrt[d]{\pi})$. Furthermore, we know that isomorphisms (6)-(9) are valid over $K^{\prime}$. We will denote these isomorphisms by $\rho_{i}^{\prime}$.
Write $u_{2}^{0}(K):=\mathrm{k}^{\mathrm{M}}_{2}(K)$ and note that ${\rm{Res\,}}_{K^{\prime}/K}(u_{2}^{i}(K))\subseteq u_{2}^{di}(K^{\prime})$. Thus, there is a natural map
${\rm{Res\,}}_{K^{\prime}/K}:u_{2}^{i}(K)/u_{2}^{i+1}(K)\to u_{2}^{di}(K^{\prime})/u_{2}^{di+1}(K^{\prime}),$ (10)
which we claim to be an isomorphism for every $0\leq i\leq\epsilon_{K}$.
Let us prove surjectivity. This amounts to proving that (the class of) the symbol on the right hand side of $\rho^{\prime}_{di}$ is in the image of the restriction morphism. Now, since these isomorphisms are well-defined regardless of the choice of a uniformizer $\pi^{\prime}$, of the elements $\omega,\varpi\in K^{\prime}$ and of the preimages $x,y\in\mathcal{O}_{K^{\prime}}$, we may choose the uniformizer to be $\pi^{\prime}:=\sqrt[d]{\pi}$ and we may always assume that $x,y\in K$ since $K$ and $K^{\prime}$ have the same residue field.
If $i=0$, we see that
${\rm{Res\,}}_{K^{\prime}/K}(\{y,\pi\})=\{y,{\pi^{\prime}}^{d}\}=d\{y,\pi^{\prime}\}.$
And since $(d,p)=1$, we get the surjectivity in this case.
If $(i,p)=1$, then we have
${\rm{Res\,}}_{K^{\prime}/K}(\{1+\omega x,y\})=\{1+\omega x,y\},$
with $v_{K^{\prime}}(\omega)=di$, so we are done as well.
If $p|i$ with $i>0$ and $p<\epsilon_{K}$, then we have
${\rm{Res\,}}_{K^{\prime}/K}(\{1+\varpi^{p}x,\pi\})=\{1+\varpi^{p}x,{\pi^{\prime}}^{d}\}=d\{1+\varpi_{i}x,\pi^{\prime}\},$
with $v_{K^{\prime}}(\varpi)=di/p$ and we are also done since $(d,p)=1$.
If $i=\epsilon_{K}$, then $(\zeta_{p}-1)^{p}=\pi^{\epsilon_{K}}u_{0}$ for some $u_{0}\in\mathcal{O}_{K^{\prime}}^{*}$. Since $K^{\prime}/K$ is totally ramified, we may fix $u\in\mathcal{O}_{K}^{*}$ such that $\bar{u}_{0}=\bar{u}$. Then,
${\rm{Res\,}}_{K^{\prime}/K}(\{1-\pi^{\epsilon_{K}}ux,\pi\})=\{1-(\zeta_{p}-1)^{p}u_{0}^{-1}ux,{\pi^{\prime}}^{d}\}=d\{1-(\zeta_{p}-1)^{p}x,\pi^{\prime}\},$ (11)
where the last equality follows from isomorphism (9) and the fact that $\bar{x}=\bar{x}\bar{u}\bar{u}_{0}^{-1}$. Since $(d,p)=1$, we are done in this case as well.
We prove now injectivity of the restriction using corestriction, which corresponds to the norm morphism $N_{K^{\prime}/K}:\mathrm{k}^{\mathrm{M}}_{2}(K^{\prime})\to\mathrm{k}^{\mathrm{M}}_{2}(K)$. Since $(d,p)=1$, we know that the composite ${\rm{Cores\,}}\circ{\rm{Res\,}}$ is a bijection on every $u_{2}^{i}(K)$. In particular, we know that it defines a bijection on the quotient $u_{2}^{i}(K)/u_{2}^{i+1}(K)$. It will suffice then to prove that this map factors through $u_{2}^{di}(K^{\prime})/u_{2}^{di+1}(K^{\prime})$, which amounts to proving that $N_{K^{\prime}/K}$ sends $u_{2}^{di+1}(K^{\prime})$ to $u_{2}^{i+1}(K)$. Note that this is evident for $i=\epsilon_{K}$ since $di=\epsilon_{K^{\prime}}$ and hence $u_{2}^{di+1}(K^{\prime})=0$, so we assume henceforth that $i<\epsilon_{K}$.
Keeping our choice of $\pi^{\prime}=\sqrt[d]{\pi}$, we know by isomorphisms (7) and (8) that any element in $u_{2}^{di+1}(K^{\prime})$ can be written as a sum of symbols of the form
  • $\{1-{\pi^{\prime}}^{j}x_{j},y_{j}\}$ (if $(j,p)=1$ and $\omega=-\pi^{\prime j}$);
  • $\{1-{\pi^{\prime}}^{j}x_{j},\pi^{\prime}\}$ (if $p|j$ and $\varpi=-\pi^{\prime j/p}$);
  • and an element in $u_{2}^{di+d}(K^{\prime})$;
where $di+1\leq j\leq di+d-1$ and $x_{j},y_{j}\in\mathcal{O}_{K}$. And since we already proved that the restriction is surjective on the quotients $u_{2}^{j}(K^{\prime})/u_{2}^{j+1}(K^{\prime})$ for any $j$, we know that $u_{2}^{di+d}(K^{\prime})$ is sent to $u_{2}^{i+1}(K)$ by the corestriction, so we may focus on the remaining symbols.
Let us compute the norm of an element of the form $1-{\pi^{\prime}}^{j}x$ with $x\in\mathcal{O}_{K}$. Write $e=(j,d)$, $d=ef$ and $j=eg$. Then the conjugates of ${\pi^{\prime}}^{j}$ are precisely $\zeta_{f}^{k}{\pi^{\prime}}^{j}$ for $10\leq k\leq f-$ 1. Thus, we get
$N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x)=\left(\prod_{k=0}^{f-1}(1-\zeta_{f}^{k}{\pi^{\prime}}^{j}x)\right)^{e}=(1-{\pi^{\prime}}^{fj}x^{f})^{e}=(1-{\pi^{\prime}}^{feg}x^{f})^{e}=(1-{\pi}^{g}x^{f})^{e}.$
Consider now a symbol of the form $\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\}$ with $x\in\mathcal{O}_{K}$ and $p|j$. Then, if we fix $m$ such that $md\equiv 1\pmod{p}$, we have
$\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\}=\{1-{\pi^{\prime}}^{j}x,{\pi^{\prime}}^{md}\}=\{1-{\pi^{\prime}}^{j}x,\pi^{m}\}=m\{1-{\pi^{\prime}}^{j}x,\pi\}.$
And thus, taking corestriction we get
$N_{K^{\prime}/K}(\{1-{\pi^{\prime}}^{j}x,\pi^{\prime}\})=N_{K^{\prime}/K}(m\{1-{\pi^{\prime}}^{j}x,\pi\})=m\{N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x),\pi\}\\ =m\{(1-{\pi}^{g}x^{f})^{e},\pi\}=me\{1-{\pi}^{g}x^{f},\pi\}.$ (12)
On the other hand, consider a symbol in $u_{2}^{j}(K^{\prime})$ of the form $\{1-{\pi^{\prime}}^{j}x,y\}$ with $x,y\in\mathcal{O}_{K}$ and $(j,p)=1$. Then, taking corestriction we have
$N_{K^{\prime}/K}(\{1-{\pi^{\prime}}^{j}x,y\})=\{N_{K^{\prime}/K}(1-{\pi^{\prime}}^{j}x),y\}=\{(1-{\pi}^{g}x^{f})^{e},y\}=e\{1-{\pi}^{g}x^{f},y\}.$ (13)
Both in (12) and (13), we get a symbol in $u_{2}^{g}(K)$. And since $g=\frac{j}{e}\geq\frac{j}{d}\geq i+\frac{1}{d}>i$, we see that the symbol lies indeed in $u_{2}^{i+1}(K)$, as wished. This concludes the proof of the claim.
Having proved that (10) is an isomorphism, fix an arbitrary uniformizer $\pi\in K$. Write $\pi={\pi^{\prime}}^{d}u_{1}$ for some uniformizer $\pi^{\prime}\in\mathcal{O}_{K^{\prime}}$ and some $u_{1}\in\mathcal{O}_{K^{\prime}}^{\times}$. Then:
  • We put $\rho_{0}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{0}^{\prime}$, where $m_{d}$ denotes multiplication by $d$ and where we use the uniformizer $\pi^{\prime}$ to define $\rho_{0}^{\prime}$. Then we see that, for a lift $y\in\mathcal{O}_{K}$ of $\bar{y}\in\bar{K}$ we have
    $\rho_{0}(\bar{y})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{y,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{y,{\pi^{\prime}}^{d}\})=\{y,\pi\},$
    where the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{y,u_{1}\}\in u_{2}^{1}(K)$ by isomorphism (1) and the fact that $\mathrm{k}^{\mathrm{M}}_{2}(\bar{K})=0$.
  • For $0<i<\epsilon_{K}$ with $(i,p)=1$ and $\omega\in K$ such that $v_{K}(\omega)=i$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ\rho_{di}^{\prime}$, where we use $\omega$ as the element with $v_{K^{\prime}}(\omega)=di$. Then we see that, for lifts $x,y\in\mathcal{O}_{K}$ of $\bar{x},\bar{y}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\frac{d\bar{y}}{\bar{y}}\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\omega x,y\})=\{1+\omega x,y\}.$
  • For $0<i<\epsilon_{K}$ with $p|i$ and $\varpi\in K$ with $v_{K}(\varpi)=i/p$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{di}^{\prime}$, where we use $\pi^{\prime}$ as the uniformizer and $\varpi$ as the element with $v_{K^{\prime}}(\varpi)=di/p$. Then we see that, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}\left(\bar{x}\pmod{\bar{K}^{p}}\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{1+\varpi^{p}x,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\varpi^{p}x,{\pi^{\prime}}^{d}\})=\{1+\varpi^{p}x,\pi\},$
    where the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{1+\varpi^{p}x,u_{1}\}\in u_{2}^{i+1}(K)$ by isomorphism (3) and the fact that $\Omega_{\bar{K}}^{1}/\mathrm{Ker}(d)\subseteq\Omega_{\bar{K}}^{2}=0$.
  • For $i=\epsilon_{K}$, we put $\rho_{i}:={\rm{Res\,}}^{-1}_{K^{\prime}/K}\circ m_{d}\circ\rho_{di}^{\prime}$, where we use $\pi^{\prime}$ as the uniformizer to define $\rho^{\prime}_{di}$. Then we see that, for a lift $x\in\mathcal{O}_{K}$ of $\bar{x}\in\bar{K}$ we have
    $\rho_{i}\left([\bar{x}]\right)={\rm{Res\,}}^{-1}_{K^{\prime}/K}(d\{1+(\zeta_{p}-1)^{p}x,\pi^{\prime}\})={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+(\zeta_{p}-1)^{p}x,{\pi^{\prime}}^{d}\})\\ ={\rm{Res\,}}^{-1}_{K^{\prime}/K}(\{1+\pi^{\epsilon_{K}}ux,{\pi^{\prime}}^{d}\})=\{1+\pi^{\epsilon_{K}}ux,\pi\},$
    where $u$ is defined as in (11) and the last equality uses the fact that $\pi={\pi^{\prime}}^{d}u_{1}$ and $\{1+\pi^{\epsilon_{K}}ux,u_{1}\}\in u_{2}^{\epsilon_{K}+1}(K)$ by isomorphism (4) and the fact that $H_{p}^{2}(\bar{K})=0$.
Now that we have extended Kato’s isomorphisms, we can follow the same proof structure as for [LPS, Prop. 1.8], which we used in the case where $K$ was assumed to contain all $p$-th roots of unity. Indeed, to do so, one may settle the following three “absorption lemmas” that are analoguous to Lemmas 1.5, 1.6 and 1.7 of loc. cit. and that will be helpful to inductively reduce every Milnor $\mathrm{K}$-theory class to a single symbol:
Lemma 4.4.
Let $i\geq 1$ with $(i,p)=1$ and let $i<j<\epsilon_{K}$. Let $a=\{1-\pi^{i}x,y\}$ be a symbol in $u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$, where $\pi$ is a uniformizer in $K$ and $x,y\in\mathcal{O}_{K}$, and let $b$ be a class in $u_{2}^{j}(K)$. Then there exist a uniformizer $\pi^{\prime}\in K$ and $x^{\prime},y^{\prime}\in\mathcal{O}_{K}$ such that
$a+b=\{1-\pi^{\prime i}x^{\prime},y^{\prime}\}+c,\text{ with }c\in u_{2}^{j+1}(K).$
Lemma 4.5.
Let $i\geq 1$ with $p|i$ and let $i<j<\epsilon_{K}$. Let $a=\{1-\pi^{i}y,\pi\}$ be a symbol in $u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$, where $\pi$ is a uniformizer in $K$ and $x\in\mathcal{O}_{K}$, and let $b$ be a class in $u_{2}^{j}(K)$. Then there exist a uniformizer $\pi^{\prime}\in K$ and $y^{\prime}\in\mathcal{O}_{K}$ such that
$a+b=\{1-\pi^{\prime i}y^{\prime},\pi^{\prime}\}+c,\text{ with }c\in u_{2}^{j+1}(K).$
Lemma 4.6.
Let $0<j<\epsilon_{K}$. Let $a=\{y,\pi\}$ be a symbol in $\mathrm{k}^{\mathrm{M}}_{2}(K)\smallsetminus u_{2}^{1}(K)$, where $\pi$ is a uniformizer in $K$ and $y\in\mathcal{O}_{K}$, and let $b$ be a class in $u_{2}^{j}(K)$. Then there exist a uniformizer $\pi^{\prime}\in K$ and $y^{\prime}\in\mathcal{O}_{K}$ such that
$a+b=\{y^{\prime},\pi^{\prime}\}+c,\text{ with }c\in u_{2}^{j+1}(K).$
The proofs being similar to those of [LPS, Lem. 1.5, 1.6, 1.7], we only provide them in the Appendix.
Let us now finish the proof Theorem 4.1 in the case $\mathrm{char}(\bar{K})=p$. To do so, consider a nontrivial class $\alpha\in\mathrm{k}^{\mathrm{M}}_{2}(K)$. Denoting $u_{2}^{0}(K):=\mathrm{k}^{\mathrm{M}}_{2}(K)$ and recalling that $u_{2}^{\epsilon_{K}+1}(K)$ is trivial, we know that there exists $0\leq i\leq\epsilon_{K}$ such that $\alpha\in u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$.
If $i=\epsilon_{K}$, then isomorphism (9) already tells us that our class is a symbol since $H^{1}_{p}(\bar{K})$ is a quotient of $\Omega_{\bar{K}}^{0}=\bar{K}$ and hence every element in $H^{1}_{p}(\bar{K})$ has the form $[\bar{x}]$ for some $\bar{x}\in\bar{K}$.
If $i<\epsilon_{K}$, fix a uniformizer $\pi$ and put $\omega:=-\pi^{i}$ if $(i,p)=1$ and $\varpi:=-\pi^{i/p}$ if $p|i$. Then we may apply one of the isomorphisms (6), (7) or (8) and see that $\alpha=a+b$, where $a$ is a symbol in $u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$ (given by the corresponding isomorphism), and $b\in u_{2}^{i+1}(K)$. We may apply then inductively one of the Lemmas 4.4, 4.5 or 4.6, so that $\alpha=a+c$, with $a$ a symbol as above and $c\in u_{2}^{\lceil\epsilon_{K}\rceil}(K)$. If $\epsilon_{K}\not\in{\mathbb{Z}}$, then we are done since we know that $u_{2}^{\lceil\epsilon_{K}\rceil}(K)=0$ in this case by Proposition 4.3. Otherwise, we have two cases:
If $p|i$ (including $i=0$), then $a=\{y,\pi\}$ for some $y\in\mathcal{O}_{K}$ and some uniformizer $\pi\in K$. On the other hand, isomorphism (9) tells us that $c=\{1-\pi^{\epsilon_{K}}ux,\pi\}$ for some $x,u\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{y(1-\pi^{\epsilon_{K}}ux),\pi\}.$
If $(i,p)=1$, then $a=\{1-\pi^{i}x,y\}$ for some $x,y\in\mathcal{O}_{K}^{\times}$ and a uniformizer $\pi\in K$. In particular, for $m$ such that $im\equiv 1\pmod{p}$,
$a=\{1-\pi^{i}x,y\}=\{1-\pi^{i}x,\pi^{im}x^{m}y\}=\{1-\pi^{i}x,\pi x^{m}y\}.$
Putting then $\pi^{\prime}:=\pi x^{m}y$ and applying isomorphism (9) to this uniformizer, we see that $c=\{1-\pi^{\epsilon_{K}}uz,\pi^{\prime}\}$ for some $z,u\in\mathcal{O}_{K}$. We have then
$\alpha=a+c=\{(1-\pi^{i}x)(1-\pi^{\epsilon_{K}}uz),\pi^{\prime}\},$
which concludes the proof of Theorem 4.1.
5.   Some remarks on fields of positive characteristic
It is natural to wonder whether Theorem 4.1 holds when $K$ has positive characteristic. As we mention in Remark 4.2, this is indeed the case if $p$ is different from the characteristic of $K$. When $\mathrm{char}(K)=p$, we know by [GS17, Thm. 9.2.4, Prop. 9.2.5] that we have an isomorphism
$\Omega_{K}^{1}/(d(\Omega_{K}^{0})+(\gamma-1)\Omega_{K}^{1})\xrightarrow{\sim}{\rm{Br}}(K)[p],$
where $\gamma$ is the inverse Cartier operator, that sends the class of $xdy$ to that of the cyclic algebra $[xy,y)$. We also know that $K=\bar{K}((t))$ with $\bar{K}$ of dimension $1$, hence with a $p$-basis consisting in a single element $y\in\bar{K}$. This tells us that $y$ and $t$ form a $p$-basis for $K$ and thus every element in $\Omega_{K}^{1}$ is a sum of two terms $\lambda dy+\mu dt$, with $\lambda,\mu\in K$. However, it seems nontrivial to prove that the quotient on the left hand side allows us to reduce such a sum to a single term, which would yield the desired result.
In a somewhat surprising way, such a computation is actually quite simple when $p=2$, to the point that it works for any field $K$ of characteristic $2$ such that $[K:K^{2}]=2$.
Proposition 5.1.
Let $K$ be a field of characteristic $2$ with $[K:K^{2}]=2$. Then every element in $\mathrm{Br}(K)[2]$ is represented by a quaternion algebra.
Proof.
Let $(y,t)$ be a $2$-basis of $K$ and take a differential $\omega=\lambda dy+\mu dt\in\Omega^{1}_{K}$. Write:
$\lambda=\lambda_{0}^{2}+\lambda_{1}^{2}y+\lambda_{2}^{2}t+\lambda_{3}^{2}yt,$
$\mu=\mu_{0}^{2}+\mu_{1}^{2}y+\mu_{2}^{2}t+\mu_{3}^{2}yt,$
with $\lambda_{1},\dots,\lambda_{4},\mu_{1},\dots,\mu_{4}\in K$. If $\mu_{3}=0$, then:
$\omega$ $=(\lambda_{0}^{2}+\lambda_{1}^{2}y+\lambda_{2}^{2}t+\lambda_{3}^{2}yt)dy+(\mu_{0}^{2}+\mu_{1}^{2}y+\mu_{2}^{2}t)dt$
$=(\lambda_{0}^{2}+\lambda_{1}^{2}y+\lambda_{2}^{2}t+\lambda_{3}^{2}yt)dy+\mu_{0}^{2}dt+\mu_{1}^{2}(d(yt)-tdy)+\mu_{2}^{2}tdt$
$\equiv((\lambda_{2}^{2}-\mu_{1}^{2})t+\lambda_{3}^{2}yt)dy\mod(d(\Omega^{0}_{K})+(\gamma-1)\Omega^{1}_{K}).$
If $\mu_{3}\neq 0$, then:
$\omega$ $=(\lambda_{0}^{2}+\lambda_{1}^{2}y+\lambda_{2}^{2}t+\lambda_{3}^{2}yt)dy+(\mu_{0}^{2}+\mu_{1}^{2}y+\mu_{2}^{2}t+\mu_{3}^{2}yt)dt$
$=\lambda_{0}^{2}dy+\lambda_{1}^{2}ydy+\lambda_{2}^{2}(d(ty)-ydt)+\lambda_{3}^{2}ytdy+(\mu_{0}^{2}+\mu_{1}^{2}y+\mu_{2}^{2}t+\mu_{3}^{2}yt)dt$
$\equiv\lambda_{3}^{2}ytdy+((\mu_{1}^{2}-\lambda_{2}^{2})y+\mu_{3}^{2}yt)dt\mod(d(\Omega^{0}_{K})+(\gamma-1)\Omega^{1}_{K})$
$\equiv ytd(\lambda_{3}^{2}y+\mu_{3}^{2}t)+(\mu_{1}^{2}-\lambda_{2}^{2})\frac{yd(\lambda_{3}^{2}y+\mu_{3}^{2}t)-\lambda_{3}^{2}ydy}{\mu_{3}^{2}}\mod(d(\Omega^{0}_{K})+(\gamma-1)\Omega^{1}_{K})$
$\equiv\left(yt+(\mu_{1}^{2}-\lambda_{2}^{2})\mu_{3}^{-2}y\right)d(\lambda_{3}^{2}y+\mu_{3}^{2}t)\mod(d(\Omega^{0}_{K})+(\gamma-1)\Omega^{1}_{K}).$
In both cases, we deduce that $\omega$ corresponds to a quaternion algebra in $\mathrm{Br}(K)[2]$. ∎
6.   Homogeneous spaces
In this section we are interested in the arithmetic of homogeneous spaces for the field $K$. We prove in particular the following instance of Serre’s Conjecture II.
Theorem 6.1.
Let $K$ be a complete discretely valued field of dimension $2$. Let $G$ be a semisimple simply connected $K$-group. Then $H^{1}(K,G)=1$.
Proof of Theorem 6.1 when $\mathrm{char}(K)=0$.
Basic reductions boil down to the case of $G$ absolutely almost $K$-simple [BFT07, §0]. According to [BP95] and [BFT07], Theorem 6.1 holds for semisimple simply connected groups of types $A$, $B$, $C$, $D$ (except trialitarian $D_{4}$), $F_{4}$ and $G_{2}$. By Proposition 3.2 and [CTGP04, Thm. 1.2(ii) & (iii)], Theorem 6.1 also holds for semisimple simply connected groups of types $D_{4}$, $E_{6}$ and $E_{7}$. It remains to deal with the case of $E_{8}$. Let $G_{0}$ be the split $K$-group of type $E_{8}$. Since $G_{0}=\mathrm{Aut}(G_{0})$, it is enough to establish that $H^{1}(K,G_{0})=1$. Since $K$ has characteristic zero, [Gil19, Lem. 9.3.2] shows that it suffices to prove the following proposition. ∎
Proposition 6.2.
Let $K$ be a complete discretely valued field of characteristic 00 with residue field $\bar{K}$ of dimension $1$. Then the $2$-dimension, the $3$-dimension and the $5$-dimension of the maximal solvable extension $L$ of $K$ of degree $2^{a}3^{b}5^{c}$ are all at most $1$.
Proof.
[…]
Remark 6.3.
Extending the proof of Proposition 6.2 to any prime $\ell$, we actually proved that $\dim(K^{\mathrm{solv}})\leq 1$, where $K^{\mathrm{solv}}$ is the maximal solvable extension of $K$.
Proof of Theorem 6.1 when $\mathrm{char}(K)>0$.
When $K$ has positive characteristic, the same reductions take us to the case of the split group of type $E_{8}$, which we denote by $G_{0}$. Here, the proof of Theorem 6.1 is a bit more involved and uses Bruhat-Tits theory. We denote by $K^{\mathrm{unr}}$ (resp. $K^{\mathrm{tame}}$) the unramified closure (resp. the tamely ramified closure) of $K$. The Galois group of $K^{\mathrm{sep}}/K^{\mathrm{tame}}$ is the wild inertia group, this is a pro-$p$-group. According to [Gil19, Thm. 8.4.1.(1)], we have $H^{1}(K^{\mathrm{tame}},G_{0})=1$. Next we have $H^{1}(K^{\mathrm{tame}}/K^{\mathrm{unr}},G_{0})=1$ according to [Gil18, Prop. 3.1]. It remains then to establish that $H^{1}(K^{\mathrm{unr}}/K,G_{0})=1$ by the method of Bruhat and Tits [BT87]. We need to check that the argument extends in this case to the situation when $k$ is not perfect.
We can reason at a finite level and we shall prove that $H^{1}(L/K,G_{0})=1$ for a given finite unramified Galois extension $L/K$. We put $\Gamma:=\mathop{\rm Gal}\nolimits(L/K)$ and denote by $\mathcal{O}_{L}/\mathcal{O}_{K}$ the Galois extension of valuation rings and by $\bar{L}/\bar{K}$ the residual Galois extension.
Let $\mathcal{B}(G_{0,L})$ be the Bruhat-Tits building of $G_{0,L}$. It comes equipped with an action of $G_{0}(L)\rtimes\Gamma$ [BT84, §4.2.12]. Let $(B_{0},T_{0})$ be a Killing couple for $G_{0}$. The split $K$–torus $T_{0}$ defines an apartment $\mathcal{A}(T_{0,L})$ of $\mathcal{B}(G_{0,L})$ which is preserved by the action of $N_{G_{0}}(T_{0})(L)\rtimes\Gamma$.
We are given a Galois cocycle $z:\Gamma\to G_{0}(L)$; it defines a section
$u_{z}:\Gamma\to G(L)\rtimes\Gamma,\,\sigma\mapsto z_{\sigma}\sigma,$
of the projection map $G_{0}(L)\rtimes\Gamma\to\Gamma$. This provides an action of $\Gamma$ on $\mathcal{B}(G_{0,L})$ called the twisted action with respect to the cocycle $z$. The Bruhat-Tits fixed point theorem [BT72, §3.2] provides a point $y\in\mathcal{B}(G_{0,L})$ which is fixed by the twisted action. This point belongs to an apartment and since $G_{0}(L)$ acts transitively on the set of apartments of $\mathcal{B}(G_{0,L})$ there exists a suitable $g\in G_{0}(L)$ such that $g^{-1}\cdot y=x\in\mathcal{A}(T_{0,L})$. We observe that $\mathcal{A}(T_{0,L})$ is fixed pointwise by $\Gamma$ (for the standard action), so that $x$ is fixed under $\Gamma$. We consider the equivalent cocycle $z^{\prime}_{\sigma}=g^{-1}\,z_{\sigma}\,\sigma(g)$ and compute
$z^{\prime}_{\sigma}\cdot x$ $=z^{\prime}_{\sigma}\cdot\sigma(x)$
$=(g^{-1}\,z_{\sigma}\,\sigma(g))(\sigma(g^{-1})\cdot\sigma(y))$
$=g^{-1}\cdot((z_{\sigma}\sigma)\cdot y)$
$=g^{-1}\,.\,y\qquad\hbox{[$y$ is fixed under the twisted action]}$]
$=x.$
Thus, without loss of generality, we may assume that $z_{\sigma}\cdot x=x$ for each $\sigma\in\Gamma$. We put $P_{x}=\mathop{\rm Stab}\nolimits_{G_{0}(L)}(x)$; since $x$ is fixed by $\Gamma$, the group $P_{x}$ is preserved by the action of $\Gamma$. Let $\mathcal{P}_{x}$ be the Bruhat-Tits $\mathcal{O}_{L}$-group scheme attached to $x$. It is smooth, we have $\mathcal{P}_{x}(\mathcal{O}_{L})=P_{x}$ and we know that its special fiber $\mathcal{P}_{x}\times_{\mathcal{O}_{L}}\bar{L}$ is smooth connected [BT84, Prop. 4.6.32, 5.2.9]. According to [McN20, Prop. 4.3.7.(a)], the unipotent radical $U_{x}$ of $\mathcal{P}_{x}\times_{\mathcal{O}_{L}}\bar{L}^{\mathrm{perf}}$ is defined over $\bar{L}$ and split over $\bar{L}$. It follows that the quotient $M_{x}=(\mathcal{P}_{x}\times_{\mathcal{O}_{L}}\bar{L})/U_{x}$ is a reductive $\bar{L}$-group. An important point is that the action of $\Gamma$ on $\mathcal{P}_{x}(\mathcal{O}_{L})$ arises from a semilinear action of $\Gamma$ on the $\mathcal{O}_{L}$–scheme $\mathcal{P}_{x}$ as explained in the beginning of [Pra20, §2]. By Galois descent, the affine $\mathcal{O}_{L}$–group scheme $\mathcal{P}_{x}$ descends to an $\mathcal{O}$-group scheme named $\mathcal{P}$. Similarly, $U_{x}$ descends to a unipotent $k$-group $U$ which is $k$-split in view of [Oes84, Ch. V, §7] and $M=(\mathcal{P}\times_{\mathcal{O}_{K}}\bar{K})/U$ is reductive.
We have that $[z_{\sigma}]$ belongs to the image of $H^{1}(\Gamma,\mathcal{P}(\mathcal{O}_{L}))\to H^{1}(\Gamma,G_{0}(L))$. We claim that we have the following bijections
$H^{1}(\Gamma,\mathcal{P}(\mathcal{O}_{L}))\cong H^{1}(\Gamma,(\mathcal{P}\times_{\mathcal{O}_{K}}\bar{K})(\bar{L}))\cong H^{1}(\Gamma,M(\bar{L})).$
The left-hand side map is bijective in view of Hensel’s Lemma [SGA3, Exp. XXIV, Prop. 8.1]. The right-hand side isomorphism refers to [GMB13, Lemme 7.3]. Since $\mathrm{scd}_{p}(\bar{K})\leq\dim_{p}(\bar{K})=1$, we have $\mathrm{scd}(\bar{K})\leq 1$, so that $H^{1}(\bar{K},M)=1$ by Serre’s Conjecture I (Steinberg’s Theorem [Ser02, III, §2.3, Thm. 1’]). It follows that $H^{1}(\Gamma,\mathcal{P}(\mathcal{O}_{L}))=1$ so that $[z_{\sigma}]=1\in H^{1}(\Gamma,\mathcal{P}(\mathcal{O}_{L}))$. ∎
In view of Examples 2.3 (b) and (d), we obtain the following statement.
Corollary 6.4.
We assume that $\bar{K}$ is separably closed of characteristic exponent $p\geq 1$ and that $[\bar{K}:\bar{K}^{p}]\leq p$. Let $G$ be a semisimple simply connected algebraic group over $K$. Then $H^{1}(K,G)=1$.
Corollary 6.5.
Let $k$ be a separably closed field of characteristic exponent $p\geq 1$ satisfying $[k:k^{p}]\leq p$. Let $G$ be a semisimple simply connected algebraic $k$-group and let $C$ be a smooth projective $k$-curve which is geometrically connected. Then the localization map $H^{1}_{\mathrm{fppf}}(C,G)\to H^{1}(k(C),G)$ is onto.
Proof.
For each point $P$ of $C$, we denote by $K_{P}$ the completion of $k(C)$ with respect to the valuation defined by $P$, by $\mathcal{O}_{P}$ its valuation ring and by $\bar{K}_{P}$ its residue field. We are given a class $\gamma\in H^{1}(k(C),G)$ and want to show that it extends to $C$. In view of Harder’s Lemma [Har67, Lem. 4.1.3] (see also [GP07, Cor. A.8]), it is enough to show that the image of $\gamma$ in $H^{1}(K_{P},G)$ comes from $H^{1}(\mathcal{O}_{P},G)$ for each closed point $P$ of $C$. According to [Bou83, IX, §3.2, Thm. 2.(c)], the field $K_{P}$ is isomorphic (but not necessarily $k$-isomorphic) to $\bar{K}_{P}((t))$. Since $\bar{K}_{P}$ is separably closed and satisfies $[\bar{K}_{P}:\bar{K}_{P}^{p}]\leq p$, Corollary 6.4 shows that $H^{1}(\bar{K}_{P},G)=1$. The local condition is then trivially fulfilled, so that the proof is complete. ∎
The case of the projective line is specially interesting.
Corollary 6.6.
Let $k$ be a field of characteristic exponent $p\geq 1$ satisfying $[k:k^{p}]\leq p$. Let $G$ be a semisimple simply connected algebraic $k$-group. Then the following hold:
  1. (1)
    $H^{1}_{\mathrm{fppf}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)=H^{1}_{\mathrm{Zar}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)$ and $H^{1}(k^{\mathrm{sep}}(t),G)=1$.
  2. (2)
    $H^{1}(k^{\mathrm{sep}}[t],G)=1$.
  3. (3)
    $H^{1}(k,G)=H^{1}(k[t],G)$.
Proof.
We prove (1). Since $H^{1}(k^{\mathrm{sep}},G)=1$, the Grothendieck-Harder theorem [Gil02, Thm. 3.8.(a)] shows that $H^{1}_{\mathrm{fppf}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)=H^{1}_{\mathrm{Zar}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)$. On the other hand, Corollary 6.5 shows that the map $H^{1}_{\mathrm{fppf}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)\to H^{1}(k^{\mathrm{sep}}(t),G)$ is onto, so that the map $H^{1}_{\mathrm{Zar}}(\mathbb{P}^{1}_{k^{\mathrm{sep}}},G)\to H^{1}(k^{\mathrm{sep}}(t),G)$ is onto. Since this map factors through $H^{1}_{\mathrm{Zar}}(k^{\mathrm{sep}}(t),G)$, we see that $H^{1}(k^{\mathrm{sep}}(t),G)=1$.
We prove (2). According to [Gil02, Cor. 3.10] we have an exact sequence of pointed sets
$H^{1}_{\mathrm{Zar}}(k^{\mathrm{sep}}[t],G)\to H^{1}_{\mathrm{fppf}}(k^{\mathrm{sep}}[t],G)\to H^{1}(k^{\mathrm{sep}}(t),G)$
so that $H^{1}_{\mathrm{Zar}}(k^{\mathrm{sep}}[t],G)=H^{1}_{\mathrm{fppf}}(k^{\mathrm{sep}}[t],G)$ in view of (1). Let $T$ be a maximal $k^{\mathrm{sep}}$-split torus of $G$. The same statement provides a surjection
${\mathrm{Hom}}_{k^{\mathrm{sep}}}(T,\mathbb{G}_{\mathrm{m}})\otimes_{\mathbb{Z}}\mathrm{Pic}(k^{\mathrm{sep}}[t])\to H^{1}_{\mathrm{Zar}}(k^{\mathrm{sep}}[t],G).$
Since the ring $k^{\mathrm{sep}}[t]$ is principal, we have $\mathrm{Pic}(k^{\mathrm{sep}}[t])=0$, so that $H^{1}_{\mathrm{Zar}}(k^{\mathrm{sep}}[t],G)=1$. Thus $H^{1}_{\mathrm{fppf}}(k^{\mathrm{sep}}[t],G)=1$.
We prove (3). Raghunathan-Ramanathan’s theorem [RR84] states that
$H^{1}(k,G)=\mathrm{Ker}(H^{1}(k[t],G)\to H^{1}(k^{\mathrm{sep}}[t],G)).$
From (2), we obtain $H^{1}(k,G)=H^{1}(k[t],G)$, as desired. ∎
Finally, Theorem 6.1 permits to extend to the $E_{8}$-case the results of Kisin-Pappas on extension of torsors [KP18, §1.4] and consequently several other results.
Appendix
For completeness, we provide here the proofs of Lemmas 4.4, 4.5 and 4.6, which boil down to explicit computations using Kato’s isomorphisms (6)–(9). Similar computations can be found in [LPS, Lem. 1.5, 1.6, 1.7].
Proof of Lemma 4.4.
Since $a\in u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$, we know by isomorphism (7) that $\bar{x}\neq 0$ and $d\bar{y}\neq 0\in\Omega_{\bar{K}}^{1}$. This last fact implies that $\bar{y}\not\in\bar{K}^{p}$ and hence $\bar{y}$ generates a $p$-basis of $\bar{K}$. In particular, every element in $\Omega_{\bar{K}}^{1}$ has the form $\bar{z}\frac{d\bar{y}}{\bar{y}}$ with $\bar{z}\in\bar{K}$.
Assume first that $(j,p)=1$. Then isomorphism (7) and the observation above tell us then that $b$ may be written as
$b=\{1-\pi^{j}z,y\}+c,\text{ with }z\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
We see then that $a+b=\{1-\pi^{i}x^{\prime},y\}+c$ with $x^{\prime}:=x+\pi^{j-i}z-\pi^{j}xz$, so we are done in this case.
Assume now that $p|j$. Fix $k$ such that $ik\equiv 1\pmod{p}$, so that
$a=\{1-\pi^{i}x,y\}=\{1-\pi^{i}x,\pi^{ik}x^{k}y\}=\{1-\pi^{i}x,\pi x^{k}y\}=\{1-{\pi^{\prime}}^{i}z,\pi^{\prime}\},$
where $\pi^{\prime}:=\pi x^{k}y$ is a uniformizer (recall that $\bar{x},\bar{y}\neq 0$, so $x,y\in\mathcal{O}_{K}^{\times}$) and $z:=\pi^{i}x{\pi^{\prime}}^{-i}\in\mathcal{O}_{K}^{\times}$. Then isomorphism (8) applied to the uniformizer $\pi^{\prime}$ tells us that $b$ may be written as
$b=\{1-{\pi^{\prime}}^{j}z^{\prime},\pi^{\prime}\}+c,\text{ with }z^{\prime}\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
We see then that $a+b=\{1-{\pi^{\prime}}^{i}x^{\prime},\pi^{\prime}\}+c$ with $x^{\prime}:=z+{\pi^{\prime}}^{j-i}z^{\prime}-{\pi^{\prime}}^{j}zz^{\prime}\in\mathcal{O}_{K}^{\times}$. In order to conclude, it suffices then to note that
$\{1-{\pi^{\prime}}^{i}x^{\prime},\pi^{\prime}\}=\{1-{\pi^{\prime}}^{i}x^{\prime},\pi^{\prime-ki+1}{x^{\prime}}^{-k}\}=\{1-{\pi^{\prime}}^{i}x^{\prime},{x^{\prime}}^{-k}\},$
so that putting $y^{\prime}:={x^{\prime}}^{-k}$ we are done. ∎
Proof of Lemma 4.5.
Assume first that $j|p$. Then isomorphism (8) tells us that $b$ may be written as
$b=\{1-{\pi}^{j}z,\pi\}+c,\text{ with }z\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
We see then that $a+b=\{1-\pi^{i}y^{\prime},\pi\}+c$ with $y^{\prime}:=y+\pi^{j-i}z-\pi^{j}yz$, so we are done in this case.
Assume now that $(j,p)=1$. Since $a\in u_{2}^{i}(K)\smallsetminus u_{2}^{i+1}(K)$, we know by isomorphism (8) that $\bar{y}\not\in\bar{K}^{p}$. This implies that $\bar{y}$ generates a $p$-basis of $\bar{K}$. In particular, every element in $\Omega_{\bar{K}}^{1}$ has the form $\bar{z}\frac{d\bar{y}}{\bar{y}}$ with $\bar{z}\in\bar{K}$. Isomorphism (7) tells us then that $b$ corresponds to a differential of the form
$\bar{z}\frac{d\bar{y}}{\bar{y}}=\sum_{l=0}^{p-1}(\bar{z}_{l}^{p}\bar{y}^{l})\frac{d\bar{y}}{\bar{y}},\text{ with }\bar{z}_{l}\in\bar{K},$
and hence it may be written as
$b=\sum_{l=0}^{p-1}\{1-{\pi}^{j}z_{l}^{p}y^{l},y\}+c,\text{ with }z_{l}\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
Now, if $l\neq 0$ and $m$ is such that $ml\equiv-1\pmod{p}$, we see that
$\{1-{\pi}^{j}z_{l}^{p}y^{l},y\}=\{1-{\pi}^{j}z_{l}^{p}y^{l},{\pi}^{mj}z_{l}^{mp}y^{ml+1}\}=\{1-{\pi}^{j}z_{l}^{p}y^{l},\pi^{mj}\}\\ =mj\{1-{\pi}^{j}z_{l}^{p}y^{l},\pi\}=\{(1-{\pi}^{j}z_{l}^{p}y^{l})^{mj},\pi\}=\{1-{\pi}^{j}w_{l},\pi\},$
for some $w_{l}\in\mathcal{O}_{K}$. Each of these terms can be absorbed by $a$ in the exact same way as we did in the case where $j|p$. Thus, we may assume that $b=\{1-\pi^{j}z_{0}^{p},y\}+c$.
Now, put $\pi^{\prime}:=\pi y$. Since $p|i$, on one hand we have
$a=\{1-\pi^{i}y,\pi\}=\{1-\pi^{i}y,\pi^{i+1}y\}=\{1-\pi^{i}y,\pi y\}=\{1-{\pi^{\prime}}^{i}y^{1-i},\pi^{\prime}\}.$
On the other hand, for $n$ such that $nj\equiv 1\pmod{p}$,
$\{1-\pi^{j}z_{0}^{p},y\}=\{1-\pi^{j}z_{0}^{p},\pi^{nj}z_{0}^{np}y\}=\{1-\pi^{j}z_{0}^{p},\pi y\}=\{1-{\pi^{\prime}}^{j}y^{-j}z_{0}^{p},\pi^{\prime}\}.$
We see then that $a+b=\{1+{\pi^{\prime}}^{i}y^{\prime},\pi^{\prime}\}+c$ with $y^{\prime}:=y^{1-i}+{\pi^{\prime}}^{j-i}y^{-j}z_{0}^{p}-{\pi^{\prime}}^{j}y^{1-i-j}z_{0}^{p}$, so we are done in this last case. ∎
Proof of Lemma 4.6.
Assume first that $j|p$. Then isomorphism (8) tells us that $b$ may be written as
$b=\{1-{\pi}^{j}z,\pi\}+c,\text{ with }z\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
We see then that $a+b=\{y^{\prime},\pi\}+c$ with $y^{\prime}:=y(1-{\pi}^{j}z)$, so we are done in this case.
Assume now that $(j,p)=1$. Since $a\in\mathrm{k}^{\mathrm{M}}_{2}(K)\smallsetminus u_{2}^{1}(K)$, we know by isomorphism (6) that $\bar{y}$ is nontrivial in $\mathrm{k}^{\mathrm{M}}_{1}(\bar{K})$ and hence $\bar{y}\not\in\bar{K}^{p}$. Thus, as in the last proof, we may assume that every element in $\Omega_{\bar{K}}^{1}$ has the form $\bar{z}\frac{d\bar{y}}{\bar{y}}$ with $\bar{z}\in\bar{K}$. Isomorphism (7) tells us then that $b$ can be written as
$b=\{1-\pi^{j}z,y\}+c,\text{ with }z\in\mathcal{O}_{K}\text{ and }c\in u_{2}^{j+1}(K).$
Finally, noting that
$\{1-\pi^{j}z,y\}=-\{y,1-\pi^{j}z\}=\{y,(1-\pi^{j}z)^{-1}\},$
we see that $a+b=\{y,\pi^{\prime}\}+c$ with $\pi^{\prime}:=\pi(1-{\pi}^{j}z)^{-1}$, so we are done in this case as well. ∎